Look at the table below with benefits and costs of a proposed project over 3 years. Year Benefits Costs / 1 2 125 150 250 75Look at the table below with benefits and costs of a proposed project over 3 years.



What’s the net present value of the project assuming a 5% discount rate?
Answer #1

First, calculate net benefit for each year as below:

Year 1 : Net benefit = 125-250 = -125

Year 2 : Net benefit = 150-75 = 75

Year 3 : Net benefit = 100-25 = 75

Now, calculate PV of these net cash flows using the formula below:

NPV = CF1/(1+r)1 + ... CFn/(1+r)n

Substituting given values,

NPV = -125/1.05 + 75/1.052 + 75/1.053

NPV = $13.76

Thus, correct option: (c) $13.8